OpenStudy (anonymous):

Factor: A.) (4a^3m+2)-(16^3m) b.) (8a^3b^3)-(27c^6d^6)

OpenStudy (anonymous):

\[(4a^{3m}+2)-16^{3m} or((4a)^{3m}+2)-16^{3m} , \not clear on your question\]

OpenStudy (anonymous):

not clear on your question

OpenStudy (anonymous):

The first one

OpenStudy (nirmalnema):

(4a^3m+2)-(16^3m) -8a^3m +2 =(-2a)^3m + 2

OpenStudy (anonymous):

how did you get that?

OpenStudy (nirmalnema):

by adding same terms./ then simplifying

OpenStudy (anonymous):

I`m still confused

OpenStudy (nirmalnema):

u can add or sub terms havinf same power

OpenStudy (nirmalnema):

having

OpenStudy (anonymous):

I`m sorry i just looked at the book and i wrote the problem wrong \[4a ^{3m+2} - 16a ^{3m}\]

OpenStudy (anonymous):

that is why.

OpenStudy (anonymous):

\[4a^{3m+2}-16a^{3m}=4(a^{3m}*a^2)-4^2a^{3m}=4a^{3m}(a^2-4)=4a^{3m}(a+2)(a-2)\]

OpenStudy (anonymous):

@aum, yours is wrong. So wrong.

OpenStudy (aum):

yep, just noticed my mistake.

OpenStudy (anonymous):

@chelseybell96 Post the second, use the equation tool. Not sure what you mean on the problem.